OpenStudy (anonymous):

F(x)= 0 for x<0, x for 0=1. Hence fi nd the cumulative distribution function of the random variable Y =-2lnX.

OpenStudy (amistre64):

I have yo wonder about this ... i think it might be the improper integral: \[\int_{0}^{1}\ -2ln(x)\ dx\] but I aint sure

OpenStudy (anonymous):

Hey, thanks for your reply, can you show me your reasoning behind the parameters 0 & 1? The irritating thing is I can do this question if I work out the parametes, I just don't know what they are.

OpenStudy (amistre64):

i was going on a gut feeling, but after reading up on it i see its wrong for 0 to 1; the cumulative is defined for everything between certain zscores; and the basic CDF is from -inf to z

OpenStudy (amistre64):

this looks like its going to have to be set up as a piecewise finction

OpenStudy (anonymous):

Yeah I've seen that, I've been researching this for about 5 hours now and have not found a single thing that either explains how to calculate the c.d.f without the p.d.f or that tells you how to work out a and b. I've also found zero examples using a log function.

OpenStudy (amistre64):

the pdf is defined by the piecewise at the start; you can draw up a pdf from that i believe ...

OpenStudy (amistre64):

F(x)= 0 for x<0, x for 0<x<1, 1 for x>=1

OpenStudy (amistre64):

another view from might be this:  but i gotta read a bit more to see what to do with these

OpenStudy (amistre64):

the pdf is the derivative of the cdf; which is why we can integrate up into it

OpenStudy (amistre64):

i know that ln(x) is undefined for values less than or equal to zero; and for a value of "1" it equals zero; so the only viable interval that makes sense to me is between 0 and 1. but again, I have no proof for that yet

OpenStudy (anonymous):

amistre please help meee

OpenStudy (amistre64):

"We'll learn how to find the probability density function of Y, using two different techniques, namely the distribution function technique and the change-of-variable technique. " https://onlinecourses.science.psu.edu/stat414/book/export/html/128

OpenStudy (amistre64):

if im reading this right, then since Y is increasing, we can use the DFT; \[F_y(y)=P(Y\le y)=P(-2ln(X)\le y)\] \[=P(-ln(X)\le \frac{1}{2}y)\] \[=P(ln(1/X)\le \frac{1}{2}y)\] \[=P(e^{ln(1/X)\le \frac{1}{2}y})\] \[=P(1/X\le e^{\frac{1}{2}y})\] \[=P(X\le e^{-\frac{1}{2}y})\] hmmm

OpenStudy (amistre64):

forgot to switch over my inequals sign ...

OpenStudy (amistre64):

\[F_y(y)=P(X\ge e^{-\frac{1}{2}y})\] might be better

OpenStudy (amistre64):

opps, I forgot that -2ln(X) is the flipped version of ln .... so its decreasing; gotta read up a little more :)

OpenStudy (amistre64):

same steps in the decreasing except for its the other side; so we use that in the 1-Fx version

OpenStudy (amistre64):

1) if its an increasing function for Y = uX, then we can find the Fx 2) if its decreasing, we go with the 1-Fx ...

OpenStudy (amistre64):

so, given this peicewise Fx stuff, lets see what we can do: F(x)= 0 for x<0, x for 0<x<1, 1 for x>=1 \[F(x)=0;\ x<0\] \[\Large 1-\int_{}^{e^{-\frac{1}{2}y}}0dx\] hmmm, just gotta see how they are getting the interval for integration ....

OpenStudy (anonymous):

Thanks a lot for all your help, I get it now :).

OpenStudy (amistre64):

ok, the interval for x on this is from (-inf,0) so the range is what we use for our integration interval since we are ingrating across the "y" whew!! .... im glad you do :) cause my workings here are a bit iffy to me